OpenStudy (anonymous):

Express the radical in simplest form: @zzr0ck3r

OpenStudy (anonymous):



OpenStudy (zzr0ck3r):

\[\sqrt{40}=\sqrt{4*10}=\sqrt{4}*\sqrt{10}=2*\sqrt{10}\]

OpenStudy (zzr0ck3r):

\[\sqrt{a*b}=\sqrt{a}*\sqrt{b}\]

OpenStudy (primeralph):

@GreenBeanDemphsey I don't think you need my help anymore, right?

OpenStudy (anonymous):

well I thought you was mad at me so I was not gonna bother you!

OpenStudy (anonymous):

@primeralph

OpenStudy (primeralph):

@GreenBeanDemphsey I don't get mad at people. I don't know how to. I simply had other people to attend to.